(1) With a speed of 0,15m/s moves a 2,4dm long conductor perpendicular to the field lines in a field flow density is 28mT. how much voltage is induced in the conductor?
(2) a car runs at a speed of 90km / h on a horizontal wall where the earth magnetic field is 52µT and the angle of inclination is 68. how much voltage is induced between the ends of a 1.6m long axis?
(3) Two parallel metal plates are in a vacuum. the voltage between the plates is 5.7kV, according to a small plastic ball with a mass of 44mg falls at a constant speed in space between the plates. the distance between the plates is 1.7cm. calculate the ball charge
1
Expert's answer
2019-05-14T10:21:23-0400

1) If the length of the conductor is 2.4 decimeters, or 24 cm, the induced voltage will be


"V=vBl\text{sin}90^\circ=0.15\cdot0.24\cdot0.028=0.001\text{ V, or 1 mV}."

2) Convert km/h to m/s:


"u=v\frac{1000}{3600}=90\frac{1000}{3600}=25\text{ m/s}."

"V=uBl\text{sin}68^\circ,"

"V=25\cdot1.6\cdot52\cdot10^{-6}\cdot0.93=1.93\cdot10^{-3}\text{ V, or 1.93 mV}."

3) The magnetic field between the plates is


"E=\frac{V}{d}."

According to Newton's second law, since the ball moves with constant speed, all forces are balanced, or mathematically:


"F_\text{electric}=F_\text{gravitational},"

"Eq=mg,"

"q=\frac{mg}{E}=\frac{mgd}{V},"

"q=\frac{44\cdot10^{-6}\cdot9.81\cdot0.017}{5700}=1.29\cdot10^{-9}\text{ C, or 1.29 nC}."